# TimeSieve: Extracting Temporal Dynamics through Information Bottleneck

### v1

@@ 3.2. IFCB
The overall architecture of the IFCB is illustrated in Figure 2.
Algorithm 1 Pseudo code of TimeSieve
0:  Time series data $X=\{x_{1},x_{2},\ldots,x_{T}\}$, true labels $Y=\{y_{T+1},y_{T+2},\ldots,y_{T+H}\}$
0:  Predicted output $\hat{Y}$
1:  Initialization: Initialize the model parameters, learning rate $\eta$, and total training epochs $E$
2:  Training:
3:  for $epoch\leftarrow 1$ to $E$ do
4:     for $i\leftarrow 1$ to $N$ do
[…]
6:        for $I\in\{A,D\}$ do
7:           $c\hat{I},\mu_{z},\Sigma_{z}\leftarrow\text{IFCB}(cI)$
8:        end for
9:        $X_{\text{reconstructed}}\leftarrow$ WRB($c\hat{A},c\hat{D}$)
10:        $\hat{Y}_{i}\leftarrow\text{MLP}(X_{\text{reconstructed}})$
11:        $\mathcal{L}_{o}\leftarrow\text{LossFunction}(\hat{Y}_{i},Y_{i})$
12:        for $I\in\{A,D\}$ do
13:           $\mathcal{L}_{IB_{I}}\leftarrow D_{KL}[\mathcal{N}(\mu_{zI},\Sigma_{zI})\,||\,\mathcal{N}(0,I)]+D_{KL}[p(z_{I})\,||\,p(z_{I}|i)]$
14:        end for
15:        $\mathcal{L}_{IB}\leftarrow\mathcal{L}_{IB_{A}}+\mathcal{L}_{IB_{D}}$
16:        $\mathcal{L}\leftarrow\mathcal{L}_{o}+\mathcal{L}_{IB}$
17:        $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
18:     end for
19:  end for
20:  Inference:
21:  for $i\leftarrow 1$ to $N$ do
[…]
23:     for $I\in\{A,D\}$ do
24:        $c\hat{I},\mu_{z},\Sigma_{z}\leftarrow\text{IFCB}(cI)$
25:     end for
26:     $X_{\text{reconstructed}}\leftarrow$ WRB($c\hat{A},c\hat{D}$)
27:     $\hat{Y}_{i}\leftarrow\text{MLP}(X_{\text{reconstructed}})$
28:  end for
29:  return  final predictions $\hat{Y}$
4. Experiments
4.1. Datasets and Baselines
We conduct comprehensive experiments to assess the performance of TimeSieve. For multivariate forecasting, we utilize a variety of real-world benchmarks, including ETT (Zhou et al., 2021), Exchange (Lai et al., 2018), Electricity(ECL) (Lin et al., 2023), and National Illness(ILI) (Zeng et al., 2022). Notably, instead of using a fixed lookback window length, we follow the approach of Koopa (Liu et al., 2023a) by setting the lookback window length $T$ to twice the forecast window length (i.e., $T=2H$). This approach is based on the fact that in real-world scenarios, historical observations are readily accessible, and utilizing a larger amount of observed data can enhance the performance of deep models, particularly as the forecast horizon extends (Liu et al., 2023a).
We extensively compare TimeSieve with state-of-the-art deep forecasting models, including Transformer-based models such as Nonstationary-Transformer (NSTformer) (Liu et al., 2023b) and Autoformer (Wu et al., 2021); MLP-based models such as LightTS (Campos et al., 2023), and Koopman operator-based models such as Koopa (Liu et al., 2023a). See the Appendix C for a more detailed introduction.
@@ 4.1. Datasets and Baselines
Table 1. Forecasting results with different forecast lengths H $\in$ $\{24,36,48,60\}$ for ILI and H $\in$ $\{48,96,144,192\}$ for others. We set the lookback length T = 2H. Red indicates the best result, while Blue refers to the second best.
@@
Models TimeSieve Koopa NSTformer LingtTS Autoformer
[…]
ETTh1 48 0.361 0.341 0.385 0.364 0.465 0.614 0.406 0.404 0.432 0.678
96 0.384 0.377 0.411 0.406 0.498 0.653 0.431 0.435 0.496 0.578
144 0.397 0.393 0.426 0.424 0.536 0.602 0.442 0.453 0.521 0.761
192 0.408 0.402 0.434 0.434 0.543 0.684 0.457 0.471 0.568 0.598
ETTh2 48 0.293 0.193 0.301 0.241 0.357 0.339 0.329 0.297 0.414 0.458
96 0.337 0.308 0.342 0.306 0.417 0.421 0.375 0.376 0.425 0.490
144 0.363 0.365 0.368 0.338 0.489 0.583 0.401 0.420 0.497 0.680
192 0.385 0.387 0.384 0.353 0.488 0.454 0.457 0.465 0.512 0.682
ETTm1 48 0.334 0.315 0.333 0.303 0.395 0.412 0.354 0.348 0.529 0.670
96 0.335 0.306 0.347 0.347 0.406 0.492 0.353 0.318 0.482 0.624
144 0.347 0.322 0.359 0.341 0.466 0.562 0.367 0.345 0.512 0.683
192 0.356 0.334 0.294 0.347 0.454 0.533 0.376 0.361 0.531 0.591
ETTm2 48 0.234 0.149 0.231 0.152 0.254 0.189 0.247 0.192 0.279 0.241
96 0.252 0.188 0.254 0.179 0.307 0.244 0.278 0.265 0.372 0.364
144 0.272 0.227 0.277 0.216 0.342 0.288 0.309 0.288 0.408 0.352
192 0.287 0.239 0.294 0.236 0.543 0.341 0.327 0.43 0.485 0.892
Exchange 48 0.139 0.045 0.149 0.046 0.187 0.073 0.159 0.067 0.205 0.124
96 0.196 0.083 0.211 0.092 0.294 0.159 0.247 0.168 0.778 0.409
144 0.244 0.113 0.265 0.141 0.375 0.292 0.272 0.31 0.68 0.671
192 0.288 0.180 0.329 0.212 0.464 0.494 0.354 0.403 0.979 0.544
Electricity 48 0.265 0.194 0.275 0.175 0.259 0.156 0.287 0.209 0.290 0.198
96 0.243 0.152 0.285 0.184 0.277 0.180 0.271 0.181 0.303 0.201
144 0.243 0.151 0.291 0.192 0.287 0.191 0.278 0.186 0.320 0.211
192 0.245 0.152 0.299 0.198 0.296 0.189 0.291 0.204 0.319 0.216
ILI 48 1.004 2.515 1.284 3.612 0.919 2.332 1.895 5.743 1.151 2.984
96 1.013 2.606 1.414 3.857 0.952 2.443 1.823 5.682 1.272 3.076
144 1.002 2.614 1.388 3.875 1.037 2.681 1.740 5.682 1.124 2.688
192 0.995 2.504 1.455 4.180 1.106 2.749 1.794 6.177 1.163 2.880
4.2. Comparative Experiment
@@
We present the forecasting results in Table 1, where Red indicates the best result and Blue refers to the second best. Across Table 1, TimeSieve consistently outperformed the baseline models in most scenarios, particularly excelling in datasets with redundant features, such as ETTh1 and ETTh2. In the ETTh1 dataset, TimeSieve achieved a significant reduction in both MAE and MSE at all prediction horizons. At the 48-time step horizon in ETTh1, TimeSieve achieved an MAE of 0.361 and an MSE of 0.341, outperforming Koopa by approximately 6.2% in MAE and 6.3% in MSE. This consistent performance can be attributed to the effective integration of the Wavelet Decomposition Block (WDB) and the Information Filtering and Compression Block (IFCB), which together capture multi-scale features while filtering out redundant features. However, it did not achieve the best results in some experiments with a forecast window of 48. We believe this may be due to the significant short-term fluctuations in time series data, making it difficult for the model to distinguish meaningful signals from redundant features in shorter forecast windows. In contrast, longer windows (e.g., 96, 144, 192) provide more historical data, enabling the model to better capture long-term dependencies and seasonal patterns, and to differentiate redundant features from short-term features.
In datasets like ETTh2 and ETTm1, the performance improvements are also notable. TimeSieve shows a considerable advantage in short to medium-term predictions (48 to 144 steps). For instance, in ETTh2 at the 48-time step horizon, TimeSieve achieved an MAE of 0.293, improving upon Koopa’s MAE of 0.301 and LightT’s 0.329.
Interestingly, the Exchange dataset showed the most significant relative improvement. At the 48-time step horizon, TimeSieve achieved an MAE of 0.139 and an MSE of 0.045, outperforming Koopa by 6.7% in MAE and 2.2% in MSE. The superior performance on this dataset can be attributed to the effective redundant features filtering by the IFCB, which is crucial for financial data often characterized by high volatility.
However, in the ILI dataset, TimeSieve’s performance, while still competitive, showed less dramatic improvements compared to other datasets. For the 48-time step horizon, TimeSieve’s MAE was 1.004, compared to Autoformer’s 1.151. The relatively modest improvement can be explained by the inherent complexity and redundant features levels in the ILI data, which present significant challenges for all models. Despite this, TimeSieve’s ability to manage and slightly outperform in such a challenging dataset.
The varying performance across different datasets and prediction horizons (48, 96, 144, 192 steps) underscores the model’s versatility. TimeSieve consistently demonstrated the substantial performance gains at shorter horizons (48 and 96 steps), where the impact of redundant features and abrupt changes is more pronounced. This is particularly evident in the ETTh1 and ETTh2 datasets, where shorter-term predictions are crucial for timely decision-making. The integrated wavelet transform and information bottleneck approach effectively balance feature extraction and redundant feature compression, ensuring a certain level of reliable predictions even when data is limited.
In conclusion, TimeSieve’s ability to achieve state-of-the-art performance on 70% of the datasets tested highlights its superior predictive accuracy and generalization capability. A comparison of model residuals for TimeSieve, Koopa, LightTS, NSTformer, and Autoformer on the ETTh2 dataset with $H=48$ and $T=96$ is shown in Figure 3, visualizing the residuals between predicted and actual values.
4.3. Ablation Study
To validate the effectiveness of the modules used in our proposed model, we conducted ablation studies focusing on the wavelet transformation and the IB components. Further discussions on the selection of wavelet basis functions and the impact of $\mathcal{L}_{IB}$ loss weights can be found in Appendix Tables 6 and Tables 7.
Wavelet Transformation.
@@ 4.3. Ablation Study
Table 2. Comparison of sequence decomposition methods with and without wavelet transformations (WDB & WRB and No Wavelet) on Exchange and ETTh2 datasets. We set the lookback length T = 2H.
@@
In Table 2, we compare the performance of our sequence decomposition methods with and without wavelet transformations(WDB&WRB and No Wavelet) on the Exchange and ETTh2 datasets. The results demonstrate that incorporating wavelet transformations significantly improves forecasting accuracy, as evidenced by lower MAE and MSE values across all forecasting horizons.
By leveraging wavelet transformations, our model can decompose the time series data into different frequency components, effectively compressing redundant features. This process enhances the model’s ability to capture the true underlying trends and patterns, leading to significant improvements in both MAE and MSE. For example, at the 96-step horizon on the Exchange dataset, the MAE decreases from 0.204 to 0.196 and the MSE decreases from 0.093 to 0.083. Similarly, on the ETTh2 dataset, the MAE decreases from 0.342 to 0.337 and the MSE decreases from 0.333 to 0.308.
@@
Table 3. Comparison of IFCB performance on Exchange and ETTh2 datasets under different conditions. $OnlyA$ applies IFCB only on the approximation coefficients after WDB, $OnlyD$ applies IFCB only on the detail coefficients after WRB, and ’TimeSieve’ applies IFCB on all coefficients.
Models TimeSieve $OnlyA$ $OnlyD$
[…]
Exchange 48 0.139 0.045 0.140 0.047 0.139 0.046
96 0.196 0.083 0.198 0.084 0.196 0.083
144 0.244 0.116 0.245 0.127 0.245 0.133
192 0.283 0.180 0.285 0.215 0.288 0.188
ETTh2 48 0.293 0.193 0.293 0.200 0.293 0.239
96 0.337 0.308 0.337 0.311 0.338 0.312
144 0.363 0.365 0.363 0.389 0.369 0.389
192 0.385 0.387 0.385 0.408 0.385 0.406
@@
In Table 3, we present a comparison of our proposed TimeSieve method against two baseline methods, $OnlyA$ and $OnlyD$, on the Exchange and ETTh2 datasets. The results indicate that TimeSieve consistently outperforms the baseline methods across all forecasting horizons. However, the improvements are more pronounced in the Exchange dataset compared to the ETTh2 dataset.
The ETTh2 dataset contains fewer redundant features, making it relatively easier for traditional methods to capture the underlying patterns effectively. Consequently, the improvements from using TimeSieve are less pronounced. By leveraging the Information Filtering and Compression Block (IFCB), TimeSieve can effectively compress redundant features while retaining critical information. This process enhances the model’s ability to capture the true underlying trends and patterns, leading to significant improvements in both MAE and MSE. For example, at the 96-step horizon, TimeSieve records a MAE of 0.337 and an MSE of 0.308, outperforming both $OnlyA$ (MAE 0.337, MSE 0.311) and $OnlyD$ (MAE 0.338, MSE 0.312). In this context, $OnlyA$ refers to applying IFCB only on the approximation coefficients, and $OnlyD$ refers to applying IFCB only on the detail coefficients.
In contrast, the Exchange dataset, being a financial exchange rate dataset, exhibits relatively more redundant features compared to ETTh2. This makes the benefit from redundant features reduction provided by IFCB more pronounced. Consequently, the improvements in MAE for TimeSieve are relatively modest compared to the baseline methods. For instance, at the 192-step horizon, TimeSieve achieves a MAE of 0.283, only slightly better than $OnlyA$ (0.285) and $OnlyD$ (0.288). However, TimeSieve demonstrates more significant improvements in MSE. Since MSE emphasizes larger errors due to squaring the error terms, it better captures the overall reduction in error magnitude provided by IFCB. At the 192-step horizon, TimeSieve reduces the MSE to 0.180, compared to 0.215 for $OnlyA$ and 0.188 for $OnlyD$. This indicates that while individual large errors (affecting MAE) may still occur, the overall error distribution is improved with TimeSieve.
The results presented in Table 3 validate the effectiveness of using IFCB for time series decomposition. By leveraging IFCB, our method can effectively capture both the coarse and fine-grained structures in the data, leading to more accurate and reliable forecasts
The results presented in Table 3 highlight the superior performance of the TimeSieve method, demonstrating the effectiveness of the Information Filtering and Compression Block (IFCB) in reducing redundant features and enhancing model accuracy.
5. Conclusions

### v2

@@ 3 TimeSieve
The overall architecture of IFCB is illustrated in Figure 2 and pseudo code of TimeSieve is shown in Algorithm 1.
@@
Table 1: Forecasting results with different forecast lengths H $\in$ $\{48,96,144,192\}$ We set the lookback length T = 2H. Red indicates the best result, while Blue refers to the second best.
@@
Models TimeSieve Koopa PatchTST DLinear NSTformer LingtTS Autoformer Metric MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE ETTh1 48 0.361 0.341 0.385 0.364 0.375 0.342 0.372 0.342 0.465 0.614 0.406 0.404 0.432 0.678 96 0.384 0.376 0.411 0.406 0.395 0.377 0.393 0.380 0.498 0.653 0.431 0.435 0.496 0.578 144 0.397 0.393 0.426 0.424 0.412 0.394 0.401 0.394 0.536 0.602 0.442 0.453 0.521 0.761 192 0.408 0.402 0.434 0.430 0.437 0.416 0.416 0.408 0.543 0.684 0.457 0.471 0.568 0.598 ETTh2 48 0.291 0.192 0.301 0.241 0.301 0.223 0.307 0.228 0.357 0.339 0.329 0.297 0.414 0.458 96 0.333 0.302 0.375 0.309 0.350 0.303 0.350 0.295 0.417 0.421 0.375 0.376 0.425 0.490 144 0.361 0.335 0.368 0.338 0.394 0.345 0.394 0.354 0.489 0.583 0.401 0.420 0.497 0.680 192 0.382 0.377 0.388 0.353 0.410 0.382 0.410 0.387 0.488 0.454 0.457 0.465 0.512 0.682 ETTm1 48 0.334 0.315 0.336 0.303 0.338 0.289 0.356 0.323 0.395 0.412 0.354 0.348 0.529 0.670 96 0.335 0.303 0.347 0.307 0.346 0.305 0.347 0.312 0.406 0.492 0.353 0.318 0.482 0.624 144 0.347 0.322 0.359 0.341 0.362 0.320 0.360 0.326 0.466 0.562 0.367 0.345 0.512 0.683 192 0.356 0.334 0.294 0.347 0.375 0.344 0.363 0.337 0.454 0.533 0.376 0.361 0.531 0.591 ETTm2 48 0.230 0.147 0.231 0.152 0.232 0.135 0.239 0.148 0.254 0.189 0.247 0.192 0.279 0.241 96 0.252 0.184 0.254 0.179 0.253 0.172 0.254 0.170 0.307 0.244 0.278 0.265 0.372 0.364 144 0.272 0.214 0.277 0.216 0.284 0.213 0.276 0.201 0.342 0.288 0.309 0.288 0.408 0.352 192 0.287 0.223 0.294 0.236 0.294 0.224 0.291 0.220 0.543 0.341 0.327 0.430 0.485 0.892 Exchange 48 0.139 0.045 0.149 0.046 0.145 0.048 0.145 0.046 0.187 0.073 0.159 0.067 0.205 0.124 96 0.195 0.083 0.211 0.092 0.204 0.090 0.223 0.089 0.294 0.159 0.247 0.168 0.778 0.409 144 0.241 0.113 0.265 0.141 0.265 0.138 0.256 0.133 0.375 0.292 0.272 0.310 0.680 0.671 192 0.284 0.180 0.329 0.212 0.298 0.181 0.301 0.182 0.464 0.494 0.354 0.403 0.979 0.544 Electricity 48 0.247 0.155 0.275 0.175 0.251 0.143 0.242 0.158 0.259 0.156 0.287 0.209 0.290 0.198 96 0.243 0.152 0.285 0.184 0.244 0.144 0.245 0.155 0.277 0.180 0.271 0.181 0.303 0.201 144 0.243 0.151 0.291 0.192 0.244 0.147 0.247 0.153 0.287 0.191 0.278 0.186 0.320 0.211 192 0.245 0.152 0.299 0.198 0.241 0.144 0.246 0.155 0.296 0.189 0.291 0.204 0.319 0.216 Weather 48 0.182 0.128 0.171 0.127 0.183 0.137 0.200 0.155 0.218 0.159 0.185 0.136 0.242 0.187 96 0.199 0.154 0.203 0.155 0.203 0.158 0.223 0.180 0.222 0.161 0.206 0.159 0.283 0.230 144 0.224 0.166 0.225 0.171 0.226 0.175 0.241 0.192 0.238 0.178 0.227 0.172 0.310 0.307 192 0.240 0.192 0.241 0.193 0.241 0.197 0.257 0.211 0.254 0.199 0.243 0.196 0.355 0.344
4 Experiments
4.1 Datasets and Baselines
We conduct comprehensive experiments to assess the performance of TimeSieve. For multivariate forecasting, we utilize a variety of real-world benchmarks, including ETT [16], Exchange [26], Electricity [27], and Weather [14]. Notably, instead of using a fixed lookback window length, we follow the approach of Koopa [28] by setting the lookback window length $T$ to twice the forecast window length (i.e., $T=2H$). This approach is based on the fact that in real-world scenarios, historical observations are readily accessible, and utilizing a larger amount of observed data can enhance the performance of deep models, particularly as the forecast horizon extends [28].
We extensively compare TimeSieve with state-of-the-art deep forecasting models, including Transformer-based models such as PatchTST [15], Nonstationary-Transformer (NSTformer) [6] and Autoformer [7]; MLP-based models such as LightTS [29], DLinear [14] and Koopman operator-based models such as Koopa [28]. See the Appendix Implementation Details for a more detailed introduction.
4.2 Comparative Experiment
@@
We present the forecasting results in Table 1, where Red indicates the best result and Blue refers to the second best. Across Table 1, TimeSieve consistently outperformed the baseline models in most scenarios, particularly excelling in datasets with redundant features, such as ETTh1 and ETTh2. In the ETTh1 dataset, TimeSieve achieved a significant reduction in both MAE and MSE at all prediction horizons. At the 48-time step horizon in ETTh1, TimeSieve achieved an MAE of 0.361 and an MSE of 0.341, outperforming Koopa by approximately 6.2% in MAE and 6.3% in MSE. This consistent performance can be attributed to the effective integration of the Wavelet Decomposition Block (WDB) and the Information Filtering and Compression Block (IFCB), which together capture multi-scale features while filtering out redundant features. However, it did not achieve the best results in some experiments with a forecast window of 48. We believe this may be due to the significant short-term fluctuations in time series data, making it difficult for the model to distinguish meaningful signals from redundant features in shorter forecast windows. In contrast, longer windows (e.g., 144, 192) provide more historical data, enabling the model to better capture long-term dependencies and seasonal patterns, and to differentiate redundant features from short-term features.
@@
Table 2: Comparison of sequence decomposition methods with and without wavelet transformations (WDB & WRB and No Wavelet) on Exchange and ETTh2 datasets. We set the lookback length T = 2H.
@@
ETTm1 48 0.339 0.331 0.334 0.315
96 0.342 0.341 0.335 0.303
144 0.359 0.347 0.347 0.322
192 0.360 0.352 0.356 0.334
Interestingly, the Exchange dataset showed the most significant relative improvement. At the 48-time step horizon, TimeSieve achieved an MAE of 0.139 and an MSE of 0.045, outperforming Koopa by 6.7% in MAE and 2.2% in MSE.
In analysis, we observed an interesting phenomenon where the model’s MSE performance is worse compared to MAE. This can be attributed to MSE’s sensitivity to outliers and large errors, as it squares the errors before averaging them. This sensitivity may impact performance, particularly because the model may filter out significant extreme values. This effect is especially pronounced in the ETTm2 dataset, which is characterized by a higher sampling frequency and an operational environment that may be more susceptible to extreme values. The formulas for calculating MSE and MAE are provided in the Appendix Metric.
The varying performance across different datasets and prediction horizons (48, 96, 144, 192 steps) underscores the model’s versatility. TimeSieve consistently demonstrated substantial performance gains at longer horizons (144 and 192 steps). This is particularly evident in the ETTm1 and ETTh1 datasets, where longer-term predictions are crucial for timely decision-making. For example, in the ETTh1 dataset at the 192-step horizon, TimeSieve achieved an MAE of 0.408 and an MSE of 0.402, showcasing substantial improvements over other models such as Koopa (MAE 0.434, MSE 0.430) and PatchTST (MAE 0.437, MSE 0.416). Similarly, in the ETTm1 dataset at the 192-step horizon, the model achieved an MAE of 0.356 and an MSE of 0.334, significantly outperforming competitors in terms of both metrics. The integrated wavelet transform and information bottleneck approach effectively balance feature extraction and redundant feature compression, ensuring a certain level of reliable predictions even when data is limited.
In conclusion, TimeSieve’s ability to achieve state-of-the-art performance on the majority of the datasets tested highlights its superior predictive accuracy and generalization capability.
4.3 Ablation Study
To validate the effectiveness of the modules used in our proposed model, we conducted ablation studies focusing on the wavelet transformation and the IFCB components. Further discussions on the selection of wavelet basis functions and the impact of $\mathcal{L}_{IB}$ loss weights can be found in Appendix Tables.
Wavelet Transformation.
Table 3: Comparison of IFCB performance on Exchange and ETTh2 datasets under different conditions. $OnlyA$ applies IFCB only on the approximation coefficients after WDB, $OnlyD$ applies IFCB only on the detail coefficients after WRB, and ’TimeSieve’ applies IFCB on all coefficients.
Models TimeSieve $OnlyA$ $OnlyD$
[…]
Exchange 48 0.139 0.045 0.141 0.048 0.139 0.047
96 0.195 0.083 0.198 0.084 0.195 0.083
144 0.241 0.113 0.245 0.125 0.245 0.132
192 0.284 0.180 0.288 0.215 0.290 0.188
ETTh2 48 0.291 0.192 0.293 0.200 0.293 0.239
96 0.333 0.302 0.337 0.311 0.338 0.312
144 0.361 0.335 0.363 0.389 0.369 0.389
192 0.382 0.377 0.385 0.408 0.385 0.406
Figure 3: The figure depicts the evolution of the mean (MU) and standard deviation (STD) within the information bottleneck as model iterations progress, with MSE as the metric, for a prediction length of 48 on the Exchange dataset.
Figure 4: The figure illustrates the changes in the mutual information related to the input $I(X;T)$ and the output $\beta I(T;Y)$ throughout model iterations, measured by MSE, for a prediction length of 48 on the Exchange dataset.
In Table 2, we compare the performance of our sequence decomposition methods with and without wavelet transformations(WDB&WRB and No Wavelet) on the Exchange, ETTh2 and ETTm1 datasets. The results demonstrate that incorporating wavelet transformations significantly improves forecasting accuracy, as evidenced by lower MAE and MSE values across all forecasting horizons.
By leveraging wavelet transformations, our model can decompose the time series data into different frequency components, effectively compressing redundant features. This process enhances the model’s ability to capture the true underlying trends and patterns, leading to significant improvements in both MAE and MSE. For example, at the 96-step horizon on the Exchange dataset, the MAE decreases from 0.204 to 0.195 and the MSE decreases from 0.093 to 0.083. Similarly, on the ETTh2 dataset, the MAE decreases from 0.342 to 0.337 and the MSE decreases from 0.333 to 0.308.
@@ 4.3 Ablation Study
In Table 3, we present a comparison of our proposed TimeSieve method against two baseline methods, $OnlyA$ and $OnlyD$, on the Exchange and ETTh2 datasets. Results show that TimeSieve consistently outperforms the baseline methods across all forecasting horizons. For instance, at a 96-step forecasting horizon on the ETTh1 dataset, TimeSieve achieved a MAE of 0.333 and an MSE of 0.302, which is superior to $OnlyA$ (MAE 0.337, MSE 0.311) and $OnlyD$ (MAE 0.338, MSE 0.312). However, improvements are more pronounced in the Exchange dataset. In shorter forecast horizons at Exchange, the enhancements are not as apparent, which may be due to the dataset’s richer set of influential features, making it challenging for IFCB to discern effectively at shorter time steps.
Figure 3 presents a scatter plot illustrating the evolution of the standard deviation and mean values in the information bottleneck as the model iterates. In this study, by observing the changes in the mean and standard deviation throughout the training process, we have identified a significant decreasing trend over time, with both parameters eventually converging within a more defined range. The convergence of mu suggests that the model’s latent representations are progressively approaching the central distribution of the data, which aids the model in identifying more representative and universally applicable features across the entire sample space. The reduction in std likely indicates an enhancement in the predictability and certainty of the model’s outputs.
Figure 4 presents a scatter plot of the mutual information values in the objective function as the model iterates. In this study, we observed that during the model training process, the mutual information related to the input $I(X;T)$ gradually decreases with increasing iterations, while the mutual information related to the output $I(T;Y)$ correspondingly increases, alongside a consistent reduction in MSE. This trend validates the effectiveness of the Information Bottleneck (IB) theory in practical applications, demonstrating that the model compresses input information while increasingly retaining information crucial for predicting the output. Specifically, the reduction in $I(X;T)$ reflects the model’s effective filtration of input data during the learning process. Concurrently, the increase in $I(T;Y)$ suggests an enhancement in the model’s capability to extract features closely associated with the task, thereby improving predictive accuracy.
5 Conclusions